High Resolution Laser Spectrscopy of the C̃^{2}A_{1}–X̃^{2}A_{1} Transition of CaNH_{2}

\underline{Zulfikar Morbi}, Chunfeng Zhao, P.F.Bernath, Center for Molecular Beams and Laser Chemistry, University of Waterloo, Waterloo, Ont., Canada, N2L 3G1.

The high resolution excitation laser spectrum of the C̃^{2}A_{1}– X̃^{2}A_{1} transition of CaNH_{2} was recorded with a laser ablation/supersonic molecular beam source. The molecule was made by ablating Ca metal with the second harmonic of a pulsed Nd:YAG laser followed by reaction of with 5% NH_{3} seeded in Ar(or He). Laser induced fluorescence spectra was recorded 12 cm downstream from the nozzle with a single-mode cw ring dye laser. Molecular constants for the K′_{a} = 0—K′′_{a} = 0 sub-band have been determined. The K′_{a} = 1 —K′′_{a} = 1 sub-band is currently being analysed.